OpenStudy (anonymous):

Let f(x) = 12 over the quantity of 4 x + 2 Find f(-1).

OpenStudy (austinl):

\[f(x)=\frac{12}{4x+2}\rightarrow f(-1)=\frac{12}{4(-1)+2}=?\]